What is the solubility of PbBr2(s) in mol/L in pure water? ..010 ..0015 1.2 x 10-6 ..017 0.0021
Answer #1

РЬ Br, —> Ph?1 +2 By - s 2 s Кsp = sx (2s - - 6.60 x 106 4ѕ3 6 - 60 x 106 s3 = 1.6S x 10 - 6 S = (-18х10-22 - 1.2 х 10 - - 0.